Home insert Runyambo Espoir & I ) 1 1/20/19 11:11 AM ECN-360-EMWF320A-Intermediate Microeconomics New Homework: Chapters 3, 4, and 19 Score: 0 of 1 pt Save Slide Se 33 of 33 (32 complete) HW Score: 61.09%, 20.16 of 33 pts Text Problem 19-16 E Question Help An increase in the market price of mens haircuts, from $15 per haircut to $25 per haircut, initially causes a local barbershop to have its employees work overtime to increase the number of daily haircuts provided from 35 to 40 When the $25 market price remains unchanged for several weeks and all other things remain equal as well, the barbershop hires additional employees and provides 55 haircuts per day What is the short-run price elasticity of supply?Your answer should have two decimal places)
Answer #1

Answer:

Since in long run the hair cut would 55 because price remain same for many weeks and production resources (Employees) also increased.

For Short run hair cut would be increase from 35 to 40

Percentage change in haircut= (40-35)/35=14.28%

Percentage change in Price=(25-15)/15=66.67%

So Short run price elasticity of supply=Percentage change in haircut/Percentage change in Price=14.28%/66.67%=0.2141